# Reconstructing Hand-Held Objects in 3D from Images and Videos

### v3

@@ Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from Internet videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Given a monocular RGB video, we aim to reconstruct hand-held object geometry in 3D, over time. In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we prompt a text-to-3D generative model using GPT-4(V) to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR). RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner. Experiments demonstrate that our approach achieves state-of-the-art performance on lab and Internet image/video datasets. We make our code and models available on the project website: https://janehwu.github.io/mcc-ho
@@ 1 Introduction
This paper addresses the problem of recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry. Given the tremendous advancements in object recognition and retrieval (e.g., generation), such as GPT-4(V) [51] and text-to-3D models [45, 57, 7, 43], a natural question that arises is whether reconstruction can benefit from leveraging these other branches of computer vision. Particularly for the challenging setting of reconstructing hand-held objects, such an approach may be essential to building up available training data to Internet scale. See Figure 1 and the supplementary videos.
Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics. First, we aim to solve a particular instance of a classic computer vision problem (along with recognition and segmentation): 3D reconstruction. It is also a particularly difficult setting because the hand and object need to be reconstructed from a single RGB image, where the hand is articulated and the object can be small with respect to the image size and partially occluded by the hand. For videos, hand-object trajectories need to additionally be temporally consistent. Second, in robotics, particularly manipulation, there is a great need for big data. Unlike other areas such as natural language processing and computer vision, where big data is already available on the Internet, large-scale trajectories of robots manipulating objects are not readily available. There has been a longstanding aspiration to leverage trajectories of human hands interacting with objects [66]; in terms of videos, this data is readily available on the Internet. However, the significant challenge is how to extract 3D hand-object trajectories from those videos. To this end, our objective is to develop an automated technique for taking raw Internet video and extracting 3D hand-object trajectories in a model-free manner. Note that this version of the problem is distinct from model-based approaches that assume a known object shape [55, 100] or use multiview methods to reconstruct the object in 3D in a separate step [13, 31].
Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment over time of the retrieved model with network-inferred geometry. In the first stage, our transformer-based model, MCC-Hand-Object (MCC-HO), estimates hand-held object geometry from each RGB image and 3D hand. In the second stage, Retrieval-Augmented Reconstruction (RAR) is used to improve reconstruction quality via recognition and retrieval. Specifically, the hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [45]). In the third stage, the retrieved 3D object is rigidly aligned with the input image or video using a combination of DINOv2 [52] and inferred 3D cues from MCC-HO. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [4, 2, 41]. To demonstrate the scalability of our approach, we additionally obtain pseudo ground truth 3D hands and objects for previously unlabeled videos in the 100DOH dataset [71].
@@
Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs hand-held object trajectories in three stages. In the first stage, MCC-HO is used to predict hand and object point clouds. In the second stage, a 3D model for the object is obtained using Retrieval-Augmented Reconstruction (RAR). In the third stage, the 3D object model is rigidly aligned with DINOv2 [52] image features and network-inferred geometry in a temporally consistent manner.
@@ 2 Related Work
Object reconstruction. Reconstructing objects from images is a longstanding problem in computer vision, and the bulk of research in this area has focused on scenarios where the object is fully or almost fully unoccluded in the input view(s). Structure from motion (SfM) [22, 70, 72, 76] and SLAM [3, 50, 73] remain robust solutions particularly for multiview reconstruction and camera parameter estimation. Learning-based object reconstruction methods typically represent 3D geometry as point clouds [60, 61], meshes [17, 86, 88], voxel grids [90, 83], or implicit surfaces [54, 47, 49, 75, 84, 89]. Following the release of Objaverse [10] and Objaverse-XL [11], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [40, 39, 42, 27]. The CO3Dv2 dataset [67] contains objects from 50 MS-COCO categories [38] and has also been used to train monocular object reconstruction models [89]. In practice, we find that off-the-shelf object reconstruction methods do not perform well on images of hand-object interactions due to occlusion from the hand, larger object pose diversity, etc.
@@
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [56, 37, 46, 48, 82, 98, 101, 53, 69, 59] or monocular video [30, 87, 92]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [68]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [81, 97]. We use HaMeR [56] in this work as it has achieved state-of-the-art performance on in-the-wild images.
@@
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [1, 2, 55, 14, 24, 25, 74, 78, 93, 91, 77, 62]. The estimated hands can then be used to rigidly align the object with images or videos [2, 55, 100] via inverse rendering [44, 64]. Due to the specific challenges arising from hand-object interactions, e.g. highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area. RHO [2] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings. RHOV [55] extended RHO to video data by adding, among other terms, temporal and optical flow losses. Aiming to add more physics-inspired constraints, HandyPriors [100] leverages a differentiable contact module to minimize relative sliding between frames. G-HOP [96] guides optimization with a generative hand-object prior.
@@
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [94, 23, 5, 6, 99, 16, 8, 58], RGB video [28, 95, 21, 96], or RGB-D video [85]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
@@
Hand-object datasets. Datasets containing hand-object interactions are either collected in lab environments [1, 20, 4, 34, 41], through large-scale capture [18, 19], or from Internet videos [9, 71, 2]. A number of hand-held object reconstruction models [5, 6, 8] were trained on either the HO-3D [20] or DexYCB [4] datasets. Since the combined number of object instances in these two datasets is 31, it is difficult for such approaches to generalize to unseen objects. HOI4D [41] is currently the largest available dataset with 3D hand and object labels (also used in [95]). There is also a comparably larger amount of data containing videos of hand-object interactions without 3D labels. Epic Kitchens [9] contains 55 hours of egocentric video of people in their personal kitchens. Ego4D [18] and Ego-Exo4D [19] are also egocentric datasets, and 100 Days of Hands (100DOH) [71] contains 131 days of footage collected from Internet videos. Following the release of [71], MOW [2] and RHOV [55] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
@@ 3 Method Overview
Our approach to reconstructing hand-object interactions from monocular RGB video combines feedforward network inference (MCC-HO) with object model retrieval (RAR) and subsequent rigid alignment. See Figure 2. We build on the insight that the presence of a human hand provides grounding in terms of both the 3D location and scale of the object. Thus, as in prior work [94, 8], the estimated object pose and shape is conditioned on the estimated 3D hand. Applying a hand-centric approach has several advantages. Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object. Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
Given a monocular RGB image or video, our method reconstructs hand-held objects in three stages. First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4). Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [45] through a technique we coin Retrieval-Augmented Reconstruction (Section 5). Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
@@ 4 Estimating Hand-Object Geometry
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [89] for the specific task of hand-object reconstruction. Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges. First, MCC assumes RGB-D images as input, whereas our model needs to work with RGB input, since our ultimate goal is to reconstruct from Internet videos. Second, we have to find a way to cope with the significant domain shift between the training datasets. MCC was trained on the CO3D dataset [67], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface. For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [32, 65], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is parameterized by the MANO hand model [68], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [56] is used to infer the hand geometry.
@@ 4.1 Architecture
MCC-HO has an encoder-decoder architecture that is composed of transformers [79] and outputs an implicit representation of hand-object geometry. The input images and 3D hand are passed to the encoder to compute a feature map that is a function of both image and hand embeddings. These features are then used to condition the decoder, which maps any 3D location to occupancy, color, and hand-object segmentation. During training, a set of query 3D points $Q$ are uniformly sampled within the bounds of a normalized volume. At test time, vertices of a voxel grid are passed as query points.
Encoder. The encoder takes as input a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a posed MANO hand mesh defined in camera world space with vertices $v_{h}\in\mathbb{R}^{778\times 3}$ and faces $f_{h}\in\mathbb{R}^{1538\times 3}$. The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [56]). In order to jointly encode the image and hand geometry, the hand mesh is rasterized and per-pixel 3D points for the hand are sampled. For each pixel $(i,j)$ in the image, a ray is cast from the camera aperture to the pixel center to determine the first intersection point with the hand mesh at a point $P_{0}(i,j)$ in camera world space. If an intersection exists, the 3D location of each point is barycentrically interpolated from the intersecting camera world space triangle vertices, denoted $\{v_{h,1},v_{h,2},v_{h,3}\}$, such that:
@@
and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [89], though we define $P$ differently, i.e., without an input depth image. The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [12] with $16\times 16$ patch embeddings. The point transformer $E^{XYZ}$ uses a self-attention-based patch embedding design that differentiates between seen and unseen pixel points. In particular, for points in $P$ where no ray-mesh intersection was found, a special $C$-dimensional embedding is learned in lieu of embedding 3D points. The function $f$ concatenates the output of the two transformers, each of which has dimensions $N^{enc}\times C$. $N^{enc}$ is the number of tokens used in the transformers and $C$ is the output channel size.
Decoder. The decoder takes as input the encoder output, $R$, as well as a set of query points $Q\in\mathbb{R}^{N_{q}\times 3}$ (we choose $N_{q}=1024$). These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [26]. Each query token is passed to three output heads that infer binary occupancy $\sigma$, RGB color $c$, and hand-object segmentation $m$, such that:

### v4

@@ Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from Internet videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Given a monocular RGB video, we aim to reconstruct hand-held object geometry in 3D, over time. In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we prompt a text-to-3D generative model using a VLM to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR). RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner. Experiments demonstrate that our approach achieves state-of-the-art performance on lab and Internet image/video datasets. We make our code and models available on the project website: https://janehwu.github.io/mcc-ho
@@ 1 Introduction
Recovering 3D hand-object interactions from visual data is an important problem in both computer vision and robotics. While there has been significant progress in reconstructing hands [pavlakos2023reconstructing, potamias2025wilor, ye2025predicting] and objects [liu2023zero, liu2024one, long2024wonder3d, xu2024instantmesh] separately, joint reconstruction of hand-object interactions presents unique challenges that cannot currently be solved by general reconstruction approaches. Existing image-to-3D models [liu2023zero, liu2024one, long2024wonder3d, xu2024instantmesh] are trained on data where the object is usually fully visible in the image; thus, these models are not trained to infer geometry primarily from hand/scene context. Procedural methods such as Structure from Motion (SfM) [tomasi1992shape] will also struggle because the object is mostly occluded, and thus it is difficult to obtain enough feature points to achieve accurate reconstruction. Instead, state-of-the-art approaches in hand-object reconstruction have relied on human assistance for data collection, whether it be lab datasets [chao2021dexycb, liu2022hoi4d, fan2023arctic] or Internet videos [cao2021reconstructing, patel2022learning]. Manually scanned or retrieved object models naturally improve reconstruction quality and can provide better initialization, but this paradigm is not scalable.
In this paper, we present a novel approach to recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry. See Figure 1 and the supplementary videos. Given an RGB image and estimated 3D hand, we train a transformer-based model, MCC-Hand-Object (MCC-HO), that jointly infers 3D hand and object geometry. This geometry is represented as a neural implicit surface composed of occupancy, color, and hand-object segmentation. By adapting our model architecture from an existing object reconstruction model, MCC [wu2023multiview], we make use of a learned object prior by fine-tuning an MCC model that is pretrained on the CO3Dv2 dataset [reizenstein2021common]. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [chao2021dexycb, cao2021reconstructing, liu2022hoi4d] (Table 1, Figures 4 and 5).
Given the limited size of existing 3D hand-object datasets, we additionally leverage vision-language models [gpt4v, comanici2025gemini] and 3D generative models [luma, poole2022dreamfusion, chen2024text, long2024wonder3d] in order to improve upon a domain-specific approach. In conjunction with MCC-HO, we propose Retrieval-Augmented Reconstruction (RAR), a method for automatically retrieving object models using large language/vision models to improve network-inferred geometry. Specifically, we prompt a VLM [gpt4v] to recognize the hand-held object in the image and provide a detailed text description. This description is passed to a text-to-3D generative model [luma] to obtain a 3D object, which is then rigidly aligned with the input image or video using a combination of DINOv2 [oquab2023dinov2] and inferred 3D cues from MCC-HO. We demonstrate that RAR leads to quantitative improvements in 3D reconstruction (Tables 2 and 5), as well as more realistic object geometries compared to existing image-to-3D methods (Figures 5 and 6).
Note that while RAR works well on average, it will fail sometimes if no close enough 3D match is found by the generative model. But the good news is that it does not need to work well on every video for the results of this framework to be useful as a “data engine” for robotics. There is a great deal of interest in getting hand-object interaction trajectories for imitation learning by robots [singh2024hand] and typically for this data is collected in lab settings [banerjee2024introducing, fu2025gigahands, hoque2025egodex] which does not result in sufficient diversity. With our paradigm, we can use a quick “thumbs-up, thumbs-down” check from a human observer to screen for all videos where the human-object interactions are reconstructed good enough. We have applied this approach with great success to Internet videos from the 100DOH dataset [shan2020understanding]; see the supplementary materials. While these videos will be of slightly lower quality than those in laboratory collected datasets, they have the advantage of scale and diversity and thus can constitute the pre-training phase of a robot data pipeline with fine-tuning provided by the laboratory datasets.
@@
Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs 3D hand-held object trajectories. First, MCC-HO is used to predict hand and object point clouds for each frame (Section 4). Then, a single 3D model for the object is obtained using Retrieval-Augmented Reconstruction (Section 5). The 3D object model is rigidly aligned with DINOv2 [oquab2023dinov2] image features and network-inferred geometry in a temporally consistent manner via our Viterbi algorithm (Section 6).
@@ 2 Related Work
Object reconstruction. Reconstructing objects from images is a longstanding problem in computer vision, and the bulk of research in this area has focused on scenarios where the object is fully or almost fully unoccluded in the input view(s). Structure from motion (SfM) [hartley2003multiple, scharstein2002taxonomy, szeliski2022computer, tomasi1992shape] and SLAM [castellanos1999spmap, mur2015orb, taketomi2017visual] remain robust solutions particularly for multiview reconstruction and camera parameter estimation. Learning-based object reconstruction methods typically represent 3D geometry as point clouds [qi2017pointnet, qi2017pointnet++], meshes [gkioxari2019mesh, wen2019pixel2mesh++, worchel2022multi], voxel grids [xie2019pix2vox, wang2021multi], or implicit surfaces [park2019deepsdf, mildenhall2021nerf, muller2022instant, tewari2020state, wang2021neus, wu2023multiview]. Following the release of Objaverse [deitke2023objaverse] and Objaverse-XL [deitke2024objaverse], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [liu2023zero, liu2024one, long2023wonder3d, hong2023lrm]. The CO3Dv2 dataset [reizenstein21co3d] contains objects from 50 MS-COCO categories [lin2014microsoft] and has also been used to train monocular object reconstruction models [wu2023multiview]. In practice, we find that off-the-shelf object reconstruction methods do not perform well on images of hand-object interactions due to occlusion from the hand, larger object pose diversity, etc.
@@
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [pavlakos2023reconstructing, li2022interacting, meng20223d, moon2023bringing, wang2023memahand, yu2023acr, zuo2023reconstructing, park2022handoccnet, rong2020frankmocap, prakash20243d] or monocular video [Karunratanakul_2023_CVPR, Wen_2023_CVPR, Xu_2023_CVPR]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [romero2022embodied]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [Wang_2023_CVPR, Yu_2023_CVPR]. We use HaMeR [pavlakos2023reconstructing] in this work as it has achieved state-of-the-art performance on in-the-wild images.
@@
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [bhatnagar2022behave, cao2021reconstructing, patel2022learning, fan2023arctic, hasson2020leveraging, hasson2021towards, tekin2019h+, tzionas2016capturing, yang2022artiboost, xie2022chore, tse2022collaborative, qu2023novel]. The estimated hands can then be used to rigidly align the object with images or videos [cao2021reconstructing, patel2022learning, zhang2023handypriors] via inverse rendering [loper2014opendr, ravi2020accelerating]. Due to the specific challenges arising from hand-object interactions, e.g. highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area. RHO [cao2021reconstructing] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings. RHOV [patel2022learning] extended RHO to video data by adding, among other terms, temporal and optical flow losses. Aiming to add more physics-inspired constraints, HandyPriors [zhang2023handypriors] leverages a differentiable contact module to minimize relative sliding between frames. G-HOP [ye2024g] guides optimization with a generative hand-object prior.
@@
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [ye2022s, hasson2019learning, chen2022alignsdf, chen2023gsdf, zhang2023moho, fu2023hacd, choi2023handnerf, Prakash2024HOI], RGB video [huang2022reconstructing, ye2023diffusion, hampali2023hand, ye2024g], or RGB-D video [wen2023bundlesdf]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
@@
Hand-object datasets. Datasets containing hand-object interactions are either collected in lab environments [bhatnagar2022behave, hampali2020honnotate, chao2021dexycb, kwon2021h2o, liu2022hoi4d], through large-scale capture [grauman2022ego4d, grauman2023ego], or from Internet videos [damen2018scaling, shan2020understanding, cao2021reconstructing]. A number of hand-held object reconstruction models [chen2022alignsdf, chen2023gsdf, choi2023handnerf] were trained on either the HO-3D [hampali2020honnotate] or DexYCB [chao2021dexycb] datasets. Since the combined number of object instances in these two datasets is 31, it is difficult for such approaches to generalize to unseen objects. HOI4D [liu2022hoi4d] is currently the largest available dataset with 3D hand and object labels (also used in [ye2023diffusion]). There is also a comparably larger amount of data containing videos of hand-object interactions without 3D labels. Epic Kitchens [damen2018scaling] contains 55 hours of egocentric video of people in their personal kitchens. Ego4D [grauman2022ego4d] and Ego-Exo4D [grauman2023ego] are also egocentric datasets, and 100 Days of Hands (100DOH) [shan2020understanding] contains 131 days of footage collected from Internet videos. Following the release of [shan2020understanding], MOW [cao2021reconstructing] and RHOV [patel2022learning] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
@@ 3 Method Overview
Our approach to reconstructing hand-object interactions from monocular RGB video combines feedforward network inference (MCC-HO) with object model retrieval (RAR) and subsequent rigid alignment. See Figure 2. First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4). Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [luma] through a technique we coin Retrieval-Augmented Reconstruction (Section 5). Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
@@ 4 Estimating Hand-Object Geometry
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [wu2023multiview] for the specific task of hand-object reconstruction. Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges. First, MCC assumes RGB-D images as input, whereas our model needs to work with RGB input, since our ultimate goal is to reconstruct from Internet videos. Second, we have to find a way to cope with the significant domain shift between the training datasets. MCC was trained on the CO3D dataset [reizenstein21co3d], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface. For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [kirillov2023segment, ravi2024sam], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is parameterized by the MANO hand model [romero2022embodied], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [pavlakos2023reconstructing] is used to infer the hand geometry.
@@ 4.1 Architecture
MCC-HO has an encoder-decoder architecture that is composed of transformers [vaswani2017attention] and outputs an implicit representation of hand-object geometry. The input images and 3D hand are passed to the encoder to compute a feature map that is a function of both image and hand embeddings. These features are then used to condition the decoder, which maps any 3D location to occupancy, color, and hand-object segmentation. During training, a set of query 3D points $Q$ are uniformly sampled within the bounds of a normalized volume. At test time, vertices of a voxel grid are passed as query points.
Encoder. The encoder takes as input a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a posed MANO hand mesh defined in camera world space with vertices $v_{h}\in\mathbb{R}^{778\times 3}$ and faces $f_{h}\in\mathbb{R}^{1538\times 3}$. The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [pavlakos2023reconstructing]). In order to jointly encode the image and hand geometry, the hand mesh is rasterized and per-pixel 3D points for the hand are sampled. For each pixel $(i,j)$ in the image, a ray is cast from the camera aperture to the pixel center to determine the first intersection point with the hand mesh at a point $P_{0}(i,j)$ in camera world space. If an intersection exists, the 3D location of each point is barycentrically interpolated from the intersecting camera world space triangle vertices, denoted $\{v_{h,1},v_{h,2},v_{h,3}\}$, such that:
@@
and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [wu2023multiview], though we define $P$ differently, i.e., without an input depth image. The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [dosovitskiy2020image] with $16\times 16$ patch embeddings. The point transformer $E^{XYZ}$ uses a self-attention-based patch embedding design that differentiates between seen and unseen pixel points. In particular, for points in $P$ where no ray-mesh intersection was found, a special $C$-dimensional embedding is learned in lieu of embedding 3D points. The function $f$ concatenates the output of the two transformers, each of which has dimensions $N^{enc}\times C$. $N^{enc}$ is the number of tokens used in the transformers and $C$ is the output channel size.
Decoder. The decoder takes as input the encoder output, $R$, as well as a set of query points $Q\in\mathbb{R}^{N_{q}\times 3}$ (we choose $N_{q}=1024$). These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [he2022masked]. Each query token is passed to three output heads that infer binary occupancy $\sigma$, RGB color $c$, and hand-object segmentation $m$, such that: